# bibliography.bib

### v1

@@
Activation Functions Considered Harmful: Recovering Neural Network Weights through Controlled Channels
Jesse Spielman and David Oswald and Mark Ryan and Jo Van Bulck School of Computer Science
University of Birmingham
Birmingham, UK
jxs1366@student.bham.ac.uk
School of Computer Science
University of Birmingham
Birmingham, UK
d.f.oswald@bham.ac.uk
School of Computer Science
University of Birmingham
Birmingham, UK
m.d.ryan@bham.ac.uk
DistriNet
KU Leuven, Belgium
Leuven, Belgium
jo.vanbulck@cs.kuleuven.be
@@ Abstract
With high-stakes machine learning applications increasingly moving to untrusted end-user or cloud environments, safeguarding pre-trained model parameters becomes essential for protecting intellectual property and user privacy. Recent advancements in hardware-isolated enclaves, notably Intel SGX, hold the promise to secure the internal state of machine learning applications even against compromised operating systems. However, we show that privileged software adversaries can exploit input-dependent memory access patterns in common neural network activation functions to extract secret weights and biases from an SGX enclave. Our attack leverages the SGX-Step framework to obtain a noise-free, instruction-granular page-access trace. In a case study of an 11-input regression network using the Tensorflow Microlite library, we demonstrate complete recovery of all first-layer weights and biases, as well as partial recovery of parameters from deeper layers under specific conditions. Our novel attack technique requires only 20 queries per input per weight to obtain all first-layer weights and biases with an average absolute error of less than 1%, improving over prior model stealing attacks.
Additionally, a broader ecosystem analysis reveals the widespread use of activation functions with input-dependent memory access patterns in popular machine learning frameworks (either directly or via underlying math libraries). Our findings highlight the limitations of deploying confidential models in SGX enclaves and emphasise the need for stricter side-channel validation of machine learning implementations, akin to the vetting efforts applied to secure cryptographic libraries.
Index Terms:
Machine Learning Security, Side Channel Attacks, Trusted Execution, Confidential Computing, SGX
1 Introduction
@@
Training and deploying Machine Learning (ML) models, in particular neural networks, often requires a massive investment in computational resources. Hence, the threat of model stealing attacks is significant, because they allow an attacker to abscond with an ML application, causing substantial financial damage as well as security and privacy risks. Attacks could be used for industrial espionage (in the case of a network that solves a difficult problem, e.g., a sophisticated automated trading system), or to ease the creation of adversarial examples to defeat a classification system, e.g., for spam filtering or intrusion detection. An application that incurs a fee to access per interaction could also be duplicated to remove this restriction. While existing model stealing attacks show how to duplicate neural networks essentially via brute force, they scale poorly for increasingly complex networks and require detailed access to output probabilities or physical access for measurements such as power consumption.
@@
Because ML (both training and inference) requires substantial computation, there is a strong focus on performance. Further, because of the hardware requirements of ML training and inference (such as in the case of an AI model which backs an end user-accessible web interface), there is an increasing need to host ML workloads in the cloud, e.g., for scaling or to use the latest Central Processing Unit (CPU) and Graphics Processing Unit (GPU) hardware. However, this move off-premises and onto (often shared) hardware introduces a number of security concerns. To mitigate the risk, substantial work has been done using Trusted Execution Environments such as Intel Software Guard Extensions (SGX) or AMD Secure Encrypted Virtualization (SEV) to protect ML workloads running on third-party hardware [russinovich2023confidential, intel_secure_ml, kumar23, li2024coreguard, kethireddy2024secure].
@@
A core component of neural networks are activation functions, chosen by developers (alongside other architectural decisions such as the number of layers or the learning rate) to introduce non-linearity. These functions are generally computed once for each neuron based on the sum of the multiplication of all incoming inputs by their weights (plus a bias term). Often, these functions are provided by a ML framework such as Tensorflow or PyTorch, which all include routines for training, running inference, and deploying models. It is important to note that these framework-provided activation functions usually rely on basic mathematical functions such as std::max() or exp(), often provided by the system’s standard libraries (e.g., glibc or musl) and optimised for speed, not side-channel resistance. TEEs like Intel SGX often supply their own (restricted) standard library (e.g., SGX’s tlibc), which provides implementations known to work within the special enclaved environment. In summary, ML frameworks inherit their security properties from the used standard library math functions.
@@
Research into side-channel attacks (both with physical access as well as remotely exploitable side channels like timing) has led to the “hardening” of certain security sensitive algorithms, especially in cryptographic libraries. However, less attention has been paid to the side-channel security of ML libraries, even though their parameters (weights and biases) are today often akin to cryptographic secrets in sensitivity. Especially in the context of TEE-protected ML workloads, the strengthened side-channel adversary model has not been sufficiently studied.
@@
In this paper, we show how pervasive data-dependent memory access patterns in activation functions across many ML frameworks can lead to the deterministic recovery of hidden model parameters when deployed in a TEE. Particularly, we develop a novel methodology to recover partial weights and biases from the first (and deeper) layer(s) of a victim network when provided with an instruction-granular page-access trace. As a practical case-study attack, we deploy a Tensorflow Microlite model inside an SGX enclave and extract deterministic, instruction-granular page access traces using the SGX-Step [van_bulck_sgx-step_2017] framework. We demonstrate successful recovery of each weight and bias of the network to more than 5 decimal places accuracy with 55 invocations of the network, or with less than 1% average error with around 20 invocations after an initial calibration phase. Our end-to-end attack outperforms Tramèr-style model stealing attacks [tramer_stealing_2016], which require around 100 invocations for each parameter in simpler classification networks. Considering the wider ecosystem, we survey widely used ML libraries and find widespread use of secret-dependent access patterns in activation functions. Furthermore, we discuss to what extent our attack vector is applicable to other popular TEE architectures.
Contributions
Summarised, our contributions are:
@@
  • An end-to-end case study of practical memory-trace extraction and accurate weight/bias recovery from Tensorflow Microlite running inside an SGX enclave.
@@
Ethical Considerations
This work contributes to better understanding limitations of widely used TEE technology. Our attacks exploit the well-documented issue of precise monitoring of enclave memory accesses using tools like SGX-Step. All experiments were conducted on proof-of-concept implementations on our own local machines.
2 Background and Related Work
Trusted Execution Environments
@@ 2 Background and Related Work
One of most commonly used and widely studied TEEs is Intel’s SGX, which, even though discontinued for client CPUs, is widely supported on Xeon Scalable server CPUs and now marketed for use cases such as secure ML deployments [intel_secure_ml]. SGX-secured code is executed in a hardware-protected ‘enclave’ (with its memory encrypted) such that even the root user cannot gain access. SGX does not make any guarantees about protections against timing and memory-based side channels [noauthor_intel_nodate], and hence is vulnerable to cache and page fault attacks [nilsson_survey_2020]. SGX also allows secrets to be securely loaded into the enclave remotely, and subsequently be (un)sealed for local storage. This allows data of the enclave (such as a pre-trained ML model) to be also protected while at rest.
@@
Other TEEs of note include ARM’s TrustZone which brings the concept of TEE to e.g., mobile devices for isolating certain software from the user, for example, in Digital Rights Management (DRM) applications [arm_tz]. Intel’s Trust Domain Extensions (TDX[intel_tdx] and AMD’s SEV [amd_sev] are TEEs that build on application-level TEEs which provide hardware-backed isolation guarantees at the level of entire Virtual Machines. Their approach is intended to make it easier for developers to take advantage of TEE security features, because applications do not need to be ported or modified for the enclaved security model.
Though not a vulnerability per se, a consequence of SGX’s adversary model is that it allows an attacker to have precise control over the enclave’s execution. This gives rise to tools like SGX-Step [van_bulck_sgx-step_2017], which allows an attacker to perfectly single-step enclaved code by causing the enclave to be interrupted, suspended, and then resumed rapidly. When paired with a more conventional side channel that can probe memory pages to detect other accesses, a precise record of a program’s execution (or trace) with respect to those pages can be recorded.
Model Stealing Attacks
[…]
Side-Channel Attacks on ML Implementations
[…]
@@
In contrast, side-channel attacks that recover the full model with physical access to the hardware and measurements of electro-magnetic (EM) emanation or power consumption have been further developed [Hua18, Batina19, Yu20, horvath2024barracuda]. However, as these attacks require measurements on or in close proximity to the targeted chips, they are of less relevance to datacenter applications, where physical access to the underlying servers is typically closely guarded.
Hence, in this paper, we focus on software-based side-channel attacks to recover weights/biases without physical access and without access to the output probabilities.
Controlled Channel Attacks
@@
For SGX, Intel explicitly places the burden of preventing secret-dependent memory accesses for security-critical code on the developer [Wenhao17]. In practice, this means that enclave developers should avoid data-dependent patterns of the form array[x] or if(x) func(), where x depends on a secret. It is important to note that in practice, such secret-dependent access patterns may only become evident at the machine code level, as optimising compilers might e.g., opt to compile “branchless” code into code with conditional branches, and vice versa.
[…]
@@
The open-source SGX-Step framework [van_bulck_sgx-step_2017, vanbulck2023sgxstep] allows for precise single-stepping of production enclaves using privileged x86 APIC timer interrupts, such that page access patterns can be deterministically monitored for every enclave instruction, giving precise insights into the operation of a victim enclave. By conservatively under-estimating the APIC timer interval, SGX-Step results in either zero or single-steps, but always avoids multi-steps [vanbulck2023sgxstep, Table 5]. Subsequent works [VanBulck18, moghimi2020copycat, aldaya2020one, van_bulck_tale_2019, constable2023aexnotify] showed that the accessed (A) bit in the enclave code page deterministically distinguishes single-steps (enclave instruction retired; A=1) from zero-steps (A=0). Intel has acknowledged perfect single stepping in a recent paper [constable2023aexnotify], which also explains the underlying mechanism to stall the CPU pipeline using microcode assists. To date, instruction-granular page-access traces extracted with SGX-Step have been repeatedly abused to reliably exploit enclave interface vulnerabilities [van_bulck_tale_2019] or to reconstruct cryptographic key material [moghimi2020copycat, aldaya2020one]. In the following, we show that such traces are sufficient to recover weights from real-world neural network libraries running inside a production SGX enclave.
3 System and Adversary Model
@@
We adhere to Intel’s standard SGX threat model, where the adversary has full root access on the victim machine, as is e.g., the case when deploying an ML model on external cloud infrastructure. Notably, we do not require physical access to the machine for our attack. Privileged software adversaries can use SGX-Step (cf. Section 2) to precisely single-step a target enclave and obtain instruction-granular page access traces. We discuss generalizations to other TEEs in Section 8
We assume the network’s weights and biases are confidential (sealed), while the enclave source code is known. The latter is a common assumption in SGX attacks [xu_controlled-channel_2015, van_bulck_sgx-step_2017, VanBulck17, Brasser17, shinde2016preventing] and the default case in the Intel SGX architecture and SDK. Non-standard confidential-code deployments only add “security through obscurity” by requiring an additional, out-of-scope reversing phase [puddu2024lack, yan_cache_2020].
We assume a pre-trained ML network has been securely loaded into an SGX enclave to perform inference. We do not consider the security of the model training process, which is orthogonal to our attacks, but believe it to be an interesting avenue for future research. Inputs are passed into the enclave and the output is returned outside the enclave through the standard ecall interface. Based on the security guarantees of SGX, an ML developer would assume that their network is protected against duplication, as the inference is handled entirely within an enclave, and weights are never stored outside.
@@
We focus on Feedforward Neural Networks because they effectively illustrate the problem we exploit and are easily deployable in SGX enclaves (e.g., without dynamic linking or requiring libc) using Tensorflow Microlite. We also assume the network architecture (at least the number of neurons per layer and activation functions) is evident from the source code, known (as in the case of a well-known architecture or a transfer-learning model), or can be recovered via a cache attack [yan_cache_2020].
4 Vulnerable Activation Functions
In this paper, we mostly consider activation functions that include a call or calls to a functions in the exp() family, e.g., sigmoid(), tanh(), and expf(). This is based on our initial observation that typical implementations of exp() contain a number of input-dependent branches with different return cases. To give an example, LABEL:lst:expf_in_tlibc shows some different code paths of the implementation of expf() in SGX’s tlibc [^1] https://github.com/intel/linux-sgx/blob/7385e1/sdk/tlibc/math/e_expf.c.
1float expf(float x)
2{
3 // ...
@@ 4 Vulnerable Activation Functions
Observing the memory access/execution patterns, e.g., with SGX-Step, allows an attacker to learn which return case an input results in, and thus obtain information on the input to exp(). Note that while several variants of exp() implementations exist (both in terms of algorithms for calculation and in terms of single and double precision floats), many of them feature the same structure, where several cases are handled with different return statements.
Attacking an activation function containing a call to exp() is similar to attacking exp() directly. One must be careful to note the sign of the input (as in the case of the $-x$ in sigmoid()’s $\exp\left(-x\right)$) and search in the correct direction. In the case of functions like tanh(), which include multiple discrete calls to exp(), the code to process the memory access traces can not assume a one-to-one relationship between a call to exp() and a single neuron’s activation. Another important consideration is the expressive power of the function, which we discuss in more detail in Section 5.3.
We summarise common activation functions in Table I, and discuss their security in Section 7.2.
Despite our focus on the exp() family of functions, any function exhibiting measurable input-dependent variation in instruction counts (or memory access patterns) is likely vulnerable.
As long as the attacker can force a vulnerable function to process inputs that fall into multiple distinguishable classes, it is possible to search for the boundary points between these cases, as we discuss in detail in Section 5. In particular, relu() does not rely on a call to exp(), but to conditional logic or a call to std::max(). While relu() is ubiquitous, especially in ML-aided computer vision applications, transformers and other modern Large Language Models use a variety of exp()-based activation functions such as softmax(), tanh() and GELU() in addition to relu(). We discuss the vulnerability of relu() in Section 7.2.1.

### v2

# Triple DES AES Advanced Encryption Standard

@@
Activation Functions Considered Harmful:
Recovering Neural Network Weights through Controlled Channels
Jesse Spielman and David Oswald and Mark Ryan and Jo Van Bulck
@@ Abstract
Recent advancements in hardware-based enclaved execution environments, such as Intel SGX, aim to protect critical model parameters in high-stakes machine learning applications increasingly moving to end-user or cloud environments. However, this also introduces the risk of privileged side-channel attacks, traditionally aimed mainly at cryptographic targets.
In this paper, we develop a novel attack methodology that exploits input-dependent memory access patterns in common neural network activation functions to extract hidden model parameters. In several case studies using the SGX-Step attack framework and the Tensorflow Microlite library, we demonstrate complete recovery of first-layer weights and biases, along with partial recovery of deeper layer parameters under specific conditions. Our novel attack technique requires only 20 queries per input per weight to obtain all first-layer weights and biases, with an average absolute error of less than 1%, improving over prior model stealing attacks. Furthermore, a broader ecosystem analysis reveals the widespread use of activation functions with input-dependent memory access patterns in popular machine learning frameworks and maths libraries. Our findings highlight the limitations of deploying confidential models in SGX enclaves and emphasise the need for stricter side-channel validation of machine learning implementations, akin to the vetting efforts applied to secure cryptographic libraries.
I Introduction
@@
Training and deploying Machine Learning (ML) models, in particular Neural Networks, often requires a massive investment in computational resources. Hence, the threat of model stealing attacks is significant, because they allow an attacker to abscond with an ML application, causing substantial financial damage as well as security and privacy risks. Attacks could be used for industrial espionage (in the case of a network that solves a difficult problem, e.g., a sophisticated automated trading system), or to ease the creation of adversarial examples to defeat a classification system, e.g., for spam filtering or intrusion detection. An application that incurs a fee to access per interaction could also be duplicated to remove this restriction. While existing model stealing attacks show how to duplicate NNs essentially via brute force, they scale poorly for increasingly complex networks and require detailed access to output probabilities or physical access for measurements such as power consumption.
@@
Because ML (both training and inference) requires substantial computation, there is a strong focus on performance. Further, because of the hardware requirements of ML training and inference (such as in the case of an AI model which backs an end user-accessible web interface), there is an increasing need to host ML workloads in the cloud. However, this move onto (often shared) hardware introduces a number of security concerns. To mitigate the risk, substantial work has been done using hardware-based Trusted Execution Environments such as Intel Software Guard Extensions (SGX) or AMD Secure Encrypted Virtualization (SEV) to protect confidential ML workloads [1, 2, 3, 4, 5]. These TEEs are presumed to facilitate the secure and efficient outsourcing of critical ML computations to third-party hardware, while simultaneously preserving the confidentiality of model parameters that embody key Intellectual Property (IP).
@@
A core component of NNs are their activation functions, chosen by developers (alongside other architectural decisions such as the number of layers or the learning rate) to introduce non-linearity. These functions are generally computed once for each neuron based on the sum of the multiplication of all incoming inputs by their weights (plus a bias term). Often, these functions are provided by a ML framework such as Tensorflow or PyTorch, which include routines for training, running inference, and deploying models. It is important to note that these framework-provided activation functions usually rely on basic mathematical functions such as std::max() or exp(), which are often provided by the system’s standard libraries (e.g., glibc) and optimised for speed, not side-channel resistance. TEEs like Intel SGX often supply their own (restricted) standard library (e.g., SGX’s tlibc), which provides implementations known to work within the special enclaved environment. Thus, ML frameworks inherit some security properties from underlying standard library maths functions.
@@
Research into Side-Channel Analysis (SCA) (both with physical access as well as remotely exploitable side channels like timing) has led to the “hardening” of certain security sensitive algorithms, especially in cryptographic libraries. However, less attention has been paid to the side-channel security of ML libraries, even though their parameters (weights and biases) are today often akin to cryptographic secrets in sensitivity. Especially in the context of TEE-protected ML workloads, the strengthened side-channel adversary model has not been sufficiently studied.
In this paper, we show how pervasive data-dependent memory access patterns in activation functions across many ML frameworks can lead to the deterministic recovery of hidden model parameters when deployed in a TEE. Particularly, we develop a novel methodology to recover partial weights and biases from the first (and deeper) layer(s) of a victim network when provided with an instruction-granular page-access trace.
@@
As a practical case-study, we deploy three Tensorflow Microlite models inside SGX enclaves and extract deterministic, instruction-granular page access traces using the SGX-Step [6] framework. Our first case study demonstrates how we can use those traces to successful recovery all first layer weights and biases to more than 5 decimal places of accuracy with 55 invocations of the network, or with less than 1% average error with around 20 invocations after an initial calibration phase. Our end-to-end attack outperforms [7] model stealing attacks [7], which require around 100 invocations for each parameter in simpler classification networks. We then use our second and third case studies to explore deeper layers and much larger networks, respectfully. Considering the wider ecosystem, we survey widely used ML libraries and find widespread use of secret-dependent access patterns in activation functions. Furthermore, we discuss to what extent our attack vector applies to other TEEs.
Contributions. Summarised, our contributions are:
@@
  • Three end-to-end case studies demonstrating concretely how, by exploiting our memory-trace extraction technique, we can perform accurate weight/bias recovery from Tensorflow Microlite networks running inside SGX enclaves in a few different configurations.
@@
Ethics and Open Science. All experiments were conducted on Proof-of-Concept (PoC) implementations on our own local machines. To ensure the reproducibility of our results, and to enable future research on side-channel attacks and defenses, we release all code and data as open-source at https://github.com/heavyimage/afch_paper.
Scope. To the best of our knowledge, this is the first study to thoroughly analyse side-channel implications of deploying ML workloads in TEEs, which are increasingly being suggested by industry and academia to preserve IP and end-user privacy. In this paper, we do not claim a fully-fledged, optimised and weaponised real-world attack; instead we focus on developing a novel attack methodology to steal partial model parameters via case-study-driven evaluation. While full recovery of real-world networks may be out of reach, we show that motivated adversaries can fully automatically extract thousands of hidden, supposedly secure parameters. Following the rich tradition in cryptography, where a long line of increasingly potent side-channel attacks have highlighted the need for constant-time programming practices, we hope our work will draw attention to the security and privacy trade-offs of using TEEs for private ML and underline the importance of side-channel-resistant coding practices.
II Background and Related Work
@@
Trusted Execution Environments. One of most widely studied TEEs is Intel’s SGX, which, even though discontinued for client CPUs, is widely supported on Xeon Scalable server Central Processing Units and now marketed for use cases such as secure ML deployments [2]. SGX-secured code is executed in a hardware-protected enclave (with its memory encrypted) such that even the root user cannot gain access. SGX also allows secrets to be securely loaded into the enclave remotely, and subsequently be (un)sealed for local storage. This allows data of the enclave (such as a pre-trained ML model) to be also protected while at rest. SGX does not provide guarantees against timing or memory-based side channels [8], rendering enclaves vulnerable to cache and page fault attacks [9, 10]. Moreover, due to SGX’s privileged adversary model, such attacks can achieve high temporal resolution using tools like SGX-Step [6], which exploit timer interrupts to single-step enclave execution, effectively interleaving enclave execution with attacker-controlled code at an instruction-level granularity.
@@
Other TEEs of note include ARM’s TrustZone which brings the concept of TEE to mobile devices [11]. Intel’s Trust Domain Extensions (TDX[12] and AMD’s SEV [13] are TEEs that provide hardware-backed isolation guarantees at the level of entire Virtual Machines.
Model Stealing Attacks. The recovery of parameters from a trained neural network via “model stealing” is well documented in the literature. Such practical recovery attacks, first demonstrated by [7], enabled the duplication (stealing) of neural networks and other models [7]. In these attacks, the original network is used as an oracle to train a duplicate network with similar hyperparameters. The attack works by choosing inputs via a search of the parameter space of the network (using one of several different heuristics) and joining that input with the corresponding output probabilities which are returned. These data pairs are used as training examples for the new network. “Solving” a network in this way results in the calculation of hidden parameters that are derived from but not necessarily precise to the original source network. Furthermore, they require a large number of queries, around 100 per parameter, as well as access to the output probabilities to essentially brute-force the space of the network. While these attacks may reconstruct the overall structure and general features of the network, it remains an open question how effective they are at reproducing the specifics of the original network, e.g., when attempting to transfer adversarial examples from duplicate to original. An alternative approach by [14] (and references therein) treat the recovery of parameters as cryptanalysis, showing that the sign information for each neuron in a 1.2M parameter CIFAR10 network can be recovered in around 30 min using a 40 GB A100 Graphics Processing Unit (GPU[14].
Side-Channel Attacks on ML Implementations. An alternative approach to recover model parameters is to use a software-based (timing or cache) side channel. Most research in this direction has focused on the recovery of hyperparameters, such as the network’s architecture, but not the precise weights: [15] demonstrate a technique for recovering the hyperparameters of a victim network such as the number of layers and their activation functions using a Flush+Reload cache attack [15]. Other works in this direction have used similar cache attacks [16], execution time [17], looked at embedded devices [18], and considered equivalent approaches on GPUs [19, 20, 21]. Other uses for software-based SCA include the recovery of network inputs from floating point timing [22], membership inference [23], and the generation of adversarial examples [24, 25].
Because the targeted leakages are much smaller and hence harder to exploit, less attention has been paid to full model recovery, i.e., weights and biases, through software-based side channels: [26] show how an attacker-controlled configuration of the x86 floating point unit can be used to recover weights from a toy NN (using a custom implementation) running in an SGX enclave [26].
[27] show that, if an adversary obtains the precise timing of each network layer, in certain cases they can use timing leakages from floating point to recover the full model [27]. In particular, the authors rely on different execution timing on certain x86 CPUs when processing “subnormal” numbers. [27] do not specify whether their attack applies to real-world ML libraries, and assume that the necessary precise timing measurements “can be achieved by either analysing the cache access pattern […] or through visual inspection of power traces”. In contrast to their work, in this paper, we consider the input-dependent memory access patterns of activation functions in widely used ML libraries, showing an end-to-end attack that uses SGX-Step for reliable side-channel observations.
@@
In contrast, side-channel attacks that recover the full model with physical access to the hardware and measurements of electro-magnetic (EM) emanation or power consumption have been further developed [28, 29, 30, 31]. However, as these attacks require measurements on or in close proximity to the targeted chips, they are of less relevance to datacenter applications, where physical access to the underlying servers is typically closely guarded. Hence, in this paper, we focus on software-based side-channel attacks to recover weights/biases without physical access and without access to the output probabilities.
@@
Controlled Channel Attacks. For SGX, Intel explicitly places the burden of preventing secret-dependent memory accesses for security-critical code on the developer [32]. In practice, this means that enclave developers should avoid data-dependent patterns of the form array[x] or if(x) func(), where x depends on a secret. It is important to note that in practice, such secret-dependent access patterns may only become evident at the machine code level, as optimising compilers might e.g., opt to compile “branchless” code into code with conditional branches, and vice versa.
Apart from well-known microarchitectural cache leakage [33], which is notoriously noisy, a number of SGX-specific side channels were discovered that exploit the privileged attacker’s control over the untrusted Operating System (OS). One of the first ones being “controlled channels” as introduced by [10] [10]. This method exploits the fact that page faults within an SGX enclave are handled by the untrusted OS. A privileged adversary can, hence, temporarily unmap a page, observe whether a page fault occurs, and from this fully deterministically infer enclave memory access patterns at a 4 KiB, page-level spatial granularity. While such page-fault attacks have been proven particularly powerful, e.g., to extract text and images [10] or full cryptographic keys [34], their relatively coarse-grained spatial resolution may limit exploitability. Consider the example code snippet provided in Figure˜1, where a tight strlen loop is executed that fits entirely within a single code and data page. As subsequent accesses to the same page are cached in the processor’s Translation Lookaside Buffer (TLB), and the enclave needs both the code and data page to make forward progress, page-fault adversaries only observe the first access to a page and are not able to distinguish successive strlen loop iterations. To overcome this limitation, [35] showed that privileged adversaries may also monitor page-table attributes [35] using the ‘accessed’ and ‘dirty’ bits to count accesses to a certain page.
@@
The open-source SGX-Step framework [6, 36] allows for precise single-stepping of production enclaves using privileged x86 APIC timer interrupts, such that page access patterns can be deterministically monitored for every enclave instruction, giving precise insights into the operation of a victim enclave. By conservatively under-estimating the APIC timer interval, SGX-Step results in either zero or single-steps, but always avoids multi-steps [36, Table 5]. Subsequent works [37, 38, 39, 40, 41] showed that the accessed (A) bit in the enclave code page deterministically distinguishes single-steps (enclave instruction retired; A=1) from zero-steps (A=0). This ability to perfectly and deterministically single-step production enclaves using SGX-Step prompted Intel to develop the opt-in AEX-Notify hardware-software mitigation (cf. Section˜VII). To date, instruction-granular page-access traces extracted with SGX-Step have been repeatedly abused to reliably exploit enclave interface vulnerabilities [40] or to reconstruct cryptographic key material [38, 39]. In the following, we show that such traces are sufficient to recover weights from real-world NN libraries running inside a production SGX enclave.
III System and Adversary Model
@@
Attacker Capabilities. We adhere to Intel’s standard SGX threat model, where the adversary has full root access on the victim machine, as is e.g., the case when deploying an ML model on external cloud infrastructure. The attacker is a software-only adversary without local physical access, for instance they may gain root access remotely over the network. Such privileged software attackers can send arbitrary inputs through the enclave software interface and can further interrupt the enclave by manipulating page tables and interrupts (cf. Section˜II). For the latter, we leverage the widely-used SGX-Step [6] framework to precisely single-step a target enclave and obtain instruction-granular page access traces. Notably, while attackers may debug enclaves using a copy of the victim code to facilitate attack development, we conduct all final attacks with SGX-Step on target enclaves running in SGX production mode, i.e., without access to debug features.
While our work focuses on the Intel SGX architecture, the underlying insights and weight-recovery techniques are applicable to other TEEs. Notably, the required side-channel primitives offered by SGX-Step, i.e., precise single-stepping and page accesses, have been demonstrated on alternative VM-based TEEs like AMD SEV [42] and Intel TDX [43, 44].
@@
Target Model. We assume a pre-trained ML network. We do not consider the training process, which is orthogonal to our attack (but we believe it to be an interesting avenue for future research). We focus on Feedforward Neural Networks because they effectively illustrate the problem we exploit and are easily deployable in SGX enclaves (e.g., without dynamic linking or requiring libc) using Tensorflow Microlite. Lastly, we assume the network architecture (at least the number of neurons per layer and activation functions) is known, e.g., it is an open-source or well-known architecture or a transfer-learning model, or can be recovered via a cache attack [15].
Target Enclave. We assume that the network has been securely loaded into a production-mode SGX enclave to perform “secure” inference. The network’s weights and biases are confidential (sealed), while the enclave source code is known. The latter is a common assumption in SGX attacks [10, 6, 35, 33, 34] and the default case in the Intel SGX architecture and SDK. Non-standard confidential-code deployments only add “security through obscurity” by requiring an additional, out-of-scope reversing phase [45, 15]. Since the attacker never has access to a self-contained binary that contains the hidden parameters, an attack such as the one proposed by [46] is not possible [46]. Inputs are passed into the enclave and the output is returned outside the enclave through the standard ecall interface. Based on the security guarantees of SGX, an ML developer would assume that their network is protected against duplication, as inference is handled entirely within an enclave and weights are never stored outside.
Activation Functions. Finally, our attack requires that the victim network use a vulnerable activation function which contains at least two input-dependent branches with unique step counts such as the time they take to return. To give an example, LABEL:lst:expf_in_tlibc shows some different code paths of the implementation of expf() in SGX’s tlibc [47]. Observing the memory access/execution patterns, e.g., with SGX-Step, allows an attacker to learn which return case an input results in, and thus obtain information on the input to exp().
For this reason, we mostly consider activation functions that include a call or calls to a functions in the exp() family, e.g., sigmoid(), tanh(), and softmax(). Notably, such activation functions are widely used in transformer architectures (even though we do not consider attacks on transformers in this paper), as with sigmoid() in Deepseek V3 [48], GELU() in BERT [49] and SwiGLU() [50] in Llama [51]. softmax() was specified in the initial transformer paper [52] and is still used in many of these implementations. Note that while several variants of exp() implementations exist, they commonly feature the same structure, where multiple error cases are handled with early return statements.
@@
Attacking an activation function containing a call to exp() is similar to attacking exp() directly. One must be careful to note the sign of the input (as in the case of the x-x in sigmoid()’s $\exp\left(-x\right)$) and search in the correct direction. In the case of functions like tanh(), which can include multiple discrete calls to exp(), the code to process the memory access traces cannot assume a one-to-one relationship between a call to exp() and a single neuron’s activation. Another important consideration is the expressive power of the function, which we discuss in more detail in Section˜IV-E.
We note NNs may access the maths functions that underpin their activations from a few different sources. Whether defined by an ML framework itself or imported externally from low level standard libraries such as glibc’s libm or even dedicated high performance maths libraries like Eigen (which may be vectorised to exploit hardware acceleration features), the code may be vulnerable to our attack. We explore this with our second case study in Section˜V-B and will more generally consider how the source of these functions can affect their vulnerability to our attack in Section˜VI.
Finally, despite our focus on the exp() family of functions, any function exhibiting measurable input-dependent variation in instruction counts (or memory access patterns) is likely vulnerable. As long as the attacker can force a vulnerable function to process inputs that fall into multiple distinguishable classes, it is possible to search for the boundary points between these cases, as we discuss in detail in Section˜IV. In particular, relu() does not rely on a call to exp(), but on conditional logic or a call to std::max(). We summarise common activation functions in Table˜I and specifically discuss the vulnerability of relu() in greater detail in Section˜VI.
IV Methodology
First, we describe the overall concept for our attack (Section˜IV-A) against a NN and then consider (Section˜IV-B) different ways ML frameworks may implement activation functions. Next (Section˜IV-C), we show how to perform our attack practically using SGX-Step to generate per-neuron traces. We then explain in more detail how to attack the first layer (Section˜IV-D) and the subsequent layers (Section˜IV-E). Finally (Section˜IV-F) we estimate the performance of our attack.
IV-A Attack Concept
The starting point for our attack is the above observation (cf. Figure˜2) that, unlike in cryptographic libraries, the maths functions that underpin many ML activation functions have not been hardened to ensure they execute in constant time and without input-dependent branches and memory access patterns. A classic example in the cache timing literature is a Look-Up Tables that may be accessed in deterministic patterns based on the function inputs. We believe this work has not been undertaken in the context of ML due to performance concerns or because security is seen as less of a focus.